find the centroid of the region with the indicated boundaries y = 4 - x^2 and the x-axis


1
Expert's answer
2022-06-02T18:15:19-0400


The coordinates of the centroid of a region bouded by curves "y=f(x)" and "y=g(x)"

("f(x) \geq g(x)" ):

"\overline{x}=\frac{1}{A}\int_a^b x[f(x)-g(x)]dx,"

"\overline{y}=\frac{1}{2A}\int_a^b [(f(x))^2-(g(x))^2]dx,"

where a and b the endpoints and A the area of the region.


We have:

"A=\int_{-2}^2 (4-x^2)dx= (4x-\frac{x^3}{3})|_{-2}^2="

"= (4\cdot 2-\frac{2^3}{3})|_{-2}^2 - (4\cdot (-2)-\frac{(-2)^3}{3})|_{-2}^2="

"=8-\frac{8}{3}+8-\frac{8}{3}=16-\frac{16}{3}=16-5\frac{1}{3}=10\frac{2}{3}."


"\overline{x}=\frac{3}{32}\int_{-2}^2 x[4-x^2]dx= \frac{3}{32}\int_{-2}^2 (4x-x^3)dx="

"=\frac{3}{32}(4\frac{x^2}{2}-\frac{x^4}{4})|_{-2}^{2}=\frac{3}{32}(2x^2-\frac{x^4}{4})|_{-2}^{2}="

"=\frac{3}{32}[(2\cdot 2^2-\frac{2^4}{4})-(2\cdot (-2)^2-\frac{(-2)^4}{4})]|_{-2}^{2}=0."


"\overline{y}=\frac{3}{64}\int_{-2}^2 (4-x^2)^2dx=\frac{3}{64}\int_{-2}^2 (16-8x^2+x^4)dx="

"=\frac{3}{64}\int_{-2}^2 (16-8x^2+x^4)dx=\frac{3}{64} (16x-8\frac{x^3}{3}+\frac{x^5}{5})|_{-2}^{2}="

"=\frac{3}{64} [(16\cdot2-\frac{8\cdot 2^3}{3}+\frac{2^5}{5})-(16\cdot(-2)-\frac{8\cdot (-2)^3}{3}+\frac{(-2)^5}{5})]="

"=\frac{3}{64}[(32-\frac{64}{3}+\frac{32}{5})-(-32+\frac{64}{3}-\frac{32}{5})]="

"=\frac{3}{64}(64-42\frac{2}{3}+12\frac{4}{5})=\frac{3}{64}(21\frac{1}{3}+12\frac{4}{5})="

"=\frac{3}{64}(21\frac{5}{15}+12\frac{12}{15})=\frac{3}{64}\cdot34\frac{2}{15}="

"=\frac{3}{64}\cdot \frac{512}{15}=\frac{8}{5}=1\frac{3}{5}."


Answer: "(0, 1\frac{3}{5})."